A particular fruit's weights are normally distributed, with a mean of 334 grams and a standard deviation of 6 grams.

If you pick 36 fruits at random, then 13% of the time, their mean weight will be greater than how many grams?
Answer #1
for normal distribution z score =(X-μ)/σx
mean μ= 334
standard deviation σ= 6
sample size       =n= 36
standard error=(σ/√n)*√(N-n)/(N-1))= 6
for top 13% or 87th percentile critical value of z=1.13
therefore corresponding value=mean+z*std deviation=335.13 grams